# High-precision direct decay energy measurements of the electron-capture decay of 97Tc

### v1

@@ IV.1 Statistical approach
In order to calculate the propagation of errors in the half-life and the 68% band in the normalized distributions of released energy as shown in Fig. 5, we employed the pseudo-experiment technique. This method involves the sampling of parameters that potentially fluctuate from probability distribution functions (PDFs). For each set of parameters the half-life times and the normalized distribution of released energy are computed, thus obtaining their PDFs. The central values are obtained as medians of the PDFs. For the two-sided bounds we took the element at and of the CPDFs. The uncertainties are reported as the difference between each bound and the median value.
We took into account and as sources of uncertainty. Assuming a Gaussian distribution of , in about cases the capture from the L1 sub-shell will not be energetically possible. Thus, we split the sampled values into groups and considered two in the current study. For the first group, the capture of the L1 electron is possible (L1 allowed), while for the second, the L1 channel is energetically forbidden, and the L2 channel is energetically possible (L1 forbidden). For each group we plotted the median and the 1 band of the normalized distribution of released energy. For the distribution of we have used a uniform distribution with values between and . We consider a number of pseudo-experiments. In particular, for the normalized distributions of released energy, we constructed a uniform grid of points for . In each point the procedure described above was applied.
@@ V Conclusion and outlook
A direct high-precision gs-to-gs EC-decay -value measurement of 97Tc (9/2+)97Mo (5/2+) was performed using the PI-ICR technique at the JYFLTRAP Penning-trap mass spectrometer. A value of 324.82(21) keV was obtained and the precision was improved by a factor of around 19 compared to literature. The measurement also improved the mass excess of 97Tc , -87219.88(26) keV/c2, by a factor of 15 compared to the previously measured value. With the refined gs-to-gs value, the candidate transition of 97Tc (9/2+)97Mo∗ (320.0 keV) was validated to be energetically allowed. This also allows an unambiguous characterization of all the possible lines in the EC spectrum of this transition at a significance level of at least 2. The transition to the 320.0-keV state with a possible allowed decay type, has a small distance of 2.0(1.0) keV for the gs-to-es to the binding energy of the electrons in the allowed daughter atomic shell L1. For further confirmation of whether this lowest emitted neutrino energy is ultra-low, the excitation energy needs to be determined with higher precision and accuracy. The atomic self-consistent many-electron Dirac–Hartree–Fock–Slater method and the nuclear shell model were utilized to predict the partial decay half-lives and energy distributions of gs-to-es EC transitions in 97Tc with low values. To calculate error propagation in half-life and the 68% confidence interval for normalized energy distributions, a pseudo-experiment technique was introduced. Multiple corrections, such as exchange, overlap, shake-up, and shake-off effects, were accounted for in these predictions for modelling the spectrum shape near the endpoint.
@@ Acknowledgements.
We acknowledge the staff of the Accelerator Laboratory of University of Jyväskylä (JYFL-ACCLAB) for providing stable online beam. We thank the support by the Academy of Finland with projects No. 306980, 312544, 275389, 284516, 295207, 314733, 315179, 327629, 320062, 354589, 345869 and 354968. The support by the EU Horizon 2020 research and innovation program under grant No. 771036 (ERC CoG MAIDEN) is acknowledged. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 861198–LISA–H2020-MSCA-ITN-2019. V.A.S., O.N., S.S., J.S., and J.K. acknowledge support from the NEPTUN project (PNRR-I8/C9-CF264, Contract No. 760100/23.05.2023 of the Romanian Ministry of Research, Innovation and Digitization). The work leading to this publication was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - AY 155/2-1.References
@@
- Aker et al. (2024) M. Aker, D. Batzler, A. Beglarian, J. Behrens, J. Beisenkötter, M. Biassoni, B. Bieringer, Y. Biondi, F. Block, S. Bobien, M. Böttcher, B. Bornschein, L. Bornschein, T. S. Caldwell, M. Carminati, A. Chatrabhuti, S. Chilingaryan, B. A. Daniel, K. Debowski, M. Descher, D. D. Barrero, P. J. Doe, O. Dragoun, G. Drexlin, F. Edzards, K. Eitel, E. Ellinger, R. Engel, S. Enomoto, A. Felden, C. Fengler, C. Fiorini, J. A. Formaggio, C. Forstner, F. M. Fränkle, K. Gauda, A. S. Gavin, W. Gil, F. Glück, S. Grohmann, R. Grössle, R. Gumbsheimer, N. Gutknecht, V. Hannen, L. Hasselmann, N. Haußmann, K. Helbing, H. Henke, S. Heyns, S. Hickford, R. Hiller, D. Hillesheimer, D. Hinz, T. Höhn, A. Huber, A. Jansen, C. Karl, J. Kellerer, K. Khosonthongkee, M. Kleifges, M. Klein, J. Kohpeiß, C. Köhler, L. Köllenberger, A. Kopmann, N. Kovač, A. Kovalík, H. Krause, L. L. Cascio, T. Lasserre, J. Lauer, T. Le, O. Lebeda, B. Lehnert, G. Li, A. Lokhov, M. Machatschek, M. Mark, A. Marsteller, E. L. Martin, C. Melzer, S. Mertens, S. Mohanty, J. Mostafa, K. Müller, A. Nava, H. Neumann, S. Niemes, A. Onillon, D. S. Parno, M. Pavan, U. Pinsook, A. W. P. Poon, J. M. L. Poyato, S. Pozzi, F. Priester, J. Ráliš, S. Ramachandran, R. G. H. Robertson, C. Rodenbeck, M. Röllig, C. Röttele, M. Ryšavý, R. Sack, A. Saenz, R. Salomon, P. Schäfer, M. Schlösser, K. Schlösser, L. Schlüter, S. Schneidewind, U. Schnurr, M. Schrank, J. Schürmann, A. Schütz, A. Schwemmer, A. Schwenck, M. Šefčík, D. Siegmann, F. Simon, F. Spanier, D. Spreng, W. Sreethawong, M. Steidl, J. Štorek, X. Stribl, M. Sturm, N. Suwonjandee, N. T. Jerome, H. H. Telle, L. A. Thorne, T. Thümmler, S. Tirolf, N. Titov, I. Tkachev, K. Urban, K. Valerius, D. Vénos, C. Weinheimer, S. Welte, J. Wendel, C. Wiesinger, J. F. Wilkerson, J. Wolf, S. Wüstling, J. Wydra, W. Xu, S. Zadorozhny, and G. Zeller, “Direct neutrino-mass measurement based on 259 days of katrin data,” (2024), arXiv:2406.13516 [nucl-ex] .

### v2

@@ IV.1 Statistical approach
In order to calculate the propagation of errors in the half-life and the 68% band in the normalized distributions of released energy as shown in Fig. 5, we employed the pseudo-experiment technique. This method involves sampling of parameters that potentially fluctuate from probability distribution functions (PDFs). For each set of parameters, the half-life times and the normalized distribution of released energy are computed, thus obtaining their PDFs. The central values are obtained as medians of the PDFs. For the two-sided bounds, we took the element at and of the CPDFs. The uncertainties are reported as the difference between each bound and the median value.
We took into account and as sources of uncertainty. Assuming a Gaussian distribution of , in about cases the capture from the L1 sub-shell will not be energetically possible. Thus, we split the sampled values into groups and considered two in the current study. For the first group, the capture of the L1 electron is possible (L1 allowed), while for the second, the L1 channel is energetically forbidden, and the L2 channel is energetically possible (L1 forbidden). For each group, we plotted the median and the 1 band of the normalized distribution of released energy. For the distribution of we have used a uniform distribution with values between and . The error propagation based on was taken into account only for the allowed transition. We consider a number of pseudo-experiments. In particular, for the normalized distributions of released energy, we constructed a uniform grid of points for . In each point, the procedure described above was applied.
@@ V Conclusion and outlook
A direct high-precision gs-to-gs EC-decay -value measurement of 97Tc (9/2+)97Mo (5/2+) was performed using the PI-ICR technique at the JYFLTRAP Penning-trap mass spectrometer. A value of 324.82(21) keV was obtained and the precision was improved by a factor of around 19 compared to literature. The measurement also improved the mass excess of 97Tc , -87219.88(26) keV/c2, by a factor of 15 compared to the previously measured value. With the refined gs-to-gs value, the candidate transition of 97Tc (9/2+)97Mo∗ (320.0 keV) was validated to be energetically allowed. This also allows an unambiguous characterization of all the possible lines in the EC spectrum of this transition at a significance level of at least 2. The transition to the 320.0-keV state has a small distance of 2.0(1.0) keV for the gs-to-es to the binding energy of the electrons in the allowed daughter atomic shell L1. For further confirmation of whether this lowest emitted neutrino energy is ultra-low, the excitation energy needs to be determined with higher precision and accuracy. Also, to define the nature of the transition, allowed or forbidden, a spectroscopic measurement of the angular momentum of the final state should be performed. A highly charged ion (HCI) Penning trap will refine the gs-to-gs to a few eV Schweiger et al. (2024), and cryogenic detectors, e.g., Transition Edge Sensors (TES) or Metallic Magnetic Calorimeters (MMCs) at low temperatures, will measure the 320.0 keV excitation energy to around 10s eV, reducing uncertainty to 10s eV. Phase-space factor analysis and gamma coincidence further constrain rates, ensuring viability for sub-eV sensitivity, as with 159Dy Gastaldo (2022). The atomic self-consistent many-electron Dirac–Hartree–Fock–Slater method and the nuclear shell model were utilized to predict the partial decay half-lives and energy distributions of gs-to-es EC transitions in 97Tc with low values. To calculate error propagation in half-life and the 68% confidence interval for normalized energy distributions, a pseudo-experiment technique was introduced. Multiple corrections, such as exchange, overlap, shake-up, and shake-off effects, were accounted for in these predictions for modeling the spectrum shape near the endpoint.
@@ Acknowledgements.
We acknowledge the staff of the Accelerator Laboratory of University of Jyväskylä (JYFL-ACCLAB) for providing stable online beam. We thank the support by the Academy of Finland with projects No. 306980, 312544, 275389, 284516, 295207, 314733, 315179, 327629, 320062, 354589, 345869 and 354968. The support by the EU Horizon 2020 research and innovation program under grant No. 771036 (ERC CoG MAIDEN) is acknowledged. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 861198–LISA–H2020-MSCA-ITN-2019. V.A.S., O.N., S.S., J.S., and J.K. acknowledge support from the NEPTUN project (PNRR-I8/C9-CF264, Contract No. 760100/23.05.2023 of the Romanian Ministry of Research, Innovation and Digitization). The work leading to this publication was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - AY 155/2-1. This project has received funding from the European Union’s Horizon Europe Research and Innovation programme under Grant Agreement No 101057511 (EURO-LABS).References
@@
- KATRIN Collaboration et al. (2025) KATRIN Collaboration et al., Science 388, 180 (2025).
@@
- Cattadori et al. (2007) C. M. Cattadori, M. De Deo, M. Laubenstein, L. Pandola, and V. I. Tretyak, Physics of Atomic Nuclei 70, 127 (2007).
@@
- Bambynek et al. (1977) W. Bambynek, H. Behrens, M. H. Chen, B. Crasemann, M. L. Fitzpatrick, K. W. D. Ledingham, H. Genz, M. Mutterer, and R. L. Intemann, Rev. Mod. Phys. 49, 77 (1977).
- Koning, Arjan et al. (2023) Koning, Arjan, Hilaire, Stephane, and Goriely, Stephane, Eur. Phys. J. A 59, 131 (2023).
- Schweiger et al. (2024) C. Schweiger, M. Braß, V. Debierre, M. Door, H. Dorrer, C. E. Düllmann, C. Enss, P. Filianin, L. Gastaldo, Z. Harman, M. W. Haverkort, J. Herkenhoff, P. Indelicato, C. H. Keitel, K. Kromer, D. Lange, Y. N. Novikov, D. Renisch, A. Rischka, R. X. Schüssler, S. Eliseev, and K. Blaum, Nature Physics 20, 921 (2024).
- Gastaldo (2022) L. Gastaldo, Journal of Low Temperature Physics 209, 804 (2022).